# Trustworthy Conceptual Explanations for Neural Networks in Robot Decision-Making

### v1

@@ III-B Interpreting the Trustworthiness of Explanations
The thresholds for probability should be decided by the practitioners depending on how much risk they are willing to take. For instance, if an engineer is using BaTCAVe to debug a manipulator used in an assembly line, the threshould can be selected leniently as the stakes might be relatively low. In such cases, we can obtain more valid explanations. In contrast, if a legislative body is using BaTCAVe to approve a new autonomous vehicle, then the thresholds should be strict. If case 1 violates, we should try new concepts to obtain a better accuracy. If case 2 violates, we can still use explanations but they might not always be the best explanations. By obtaining the mean score, we can obtain an average explanation.
IV Experiments
Figure 3: a) JetBot b) Lift cube c) Pick-and-place d) Nut assembly
We consider three types of algorithms for learning robot decision-making: binary supervised learning, behavioral cloning (BC), and proximal policy optimization (PPO). The experimental setup (also aailable on Github), results, and findings are detailed below.
Figure 4: The effects of (a) fine-tuning and (b) data augmentation. The higher the score, the better the explanation is. (c) Shows while common XAI methods can highlight the orange box, they do not reveal what attributes of the box contribute to the decision of the DNN, making it harder for the engineers to improve the DNN based on the explanations. In contrast, BaTCAVe provides semantically meaningful explanations. (d) Highlights the change in confidence over modifying darkness factor in input with models trained with different data augmentation (C-Modification).
IV-A Experiment 1: Mobile Robot Navigation Using Vision
@@
Results: With the objective of avoiding orange obstacles, we fine-tuned the full DNN with images of an orange box, shown in Fig. 3a, until we achieve a validation accuracy of 100%. We surveyed 20 engineering students who have at least 1 year of experience training DNNs to get their opinion on what the DNN has learned just by reading our description. We described them the architecture and showed fine-tuning images without telling them that our objective is avoiding orange obstacles. First, we asked them to describe what attributes the DNN should have learned and then we gave them a list of potential concepts to narrow down their choices. Many human speculated that the model will be fined-tuned to distinguish orange from the rest.
In contract, showcasing the BaTCAVe’s ability to provide true explanations that engineers might not even think of, as shown in Fig. 4(a), BaTCAVe revealed that the model has learned to distinguish dark from light objects (or the value or brightness in HSV scale). The orange box is merely a shade of the broader “dark” concept. Similar to C-deletion in XAI literature [20, 24, 25], we verified that the darkness is an important concept by gradually varying the brightness of the orange object, as shown in Fig. 4(d). Additionally, based on Fig. 4, BaTCAVe made the following explanations:
@@
  • As shown in Fig 4(b), the importance of the dark concept remains consistent across different data augmentation methods—adding color jitter (CJ) and/or image rotation (R)—for LFT while it varies for FFT.
Since XAI methods are hard to quantitatively benchmark, similar to other work, Fig. 4(c) qualitatively compares why BaTCAVe is a better choice than feature attribution methods for robot learning.
IV-B Experiment 2: Lift Cube, Pick & Place, and Nut Assembly with Proprioceptive Sensors
TABLE I: BaTCAVe scores across different tasks along with uncertainty estimates. Each score measures the impact of $C_{I}$.
Task Action Concept object eef_pos eef_quat gripper
Lift Cube $\Delta$ X ()($\uparrow$)( ↑ ) $0.84\pm 0.31$ $0.42\pm 0.40$ $0.42\pm 0.49$ $0.42\pm 0.49$
$\Delta$ Z ()($\uparrow$)( ↑ ) $0.98\pm 0.10$ $0.99\pm 0.04$ $0.42\pm 0.49$ $0.39\pm 0.48$
Pick & place $\Delta$ X ()($\uparrow$)( ↑ ) $1\pm 0.0$ $0.54\pm 0.49$ $0.96\pm 0.16$ $0.09\pm 0.09$
$\Delta$ Z ()($\uparrow$)( ↑ ) $0.9\pm 0.17$ $0.85\pm 0.35$ $0.77\pm 0.41$ $0.91\pm 0.27$
Nut Assembly $\Delta$ X ()($\uparrow$)( ↑ ) $0.64\pm 0.47$ $0.59\pm 0.49$ $0.44\pm 0.49$ $0.87\pm 0.32$
$\Delta$ Y ()($\uparrow$)( ↑ ) $0.87\pm 0.32$ $0.59\pm 0.48$ $0.46\pm 0.49$ $0.37\pm 0.48$
Setup: We use a Panda, a 6-DOF robotic arm with a gripper, to complete the three tasks shown in Fig 3(b,c,d) on RoboSuite [31]. The setup includes various proprioceptive sensors to monitor the arm’s movements and positions. By using the DNN shown in Fig 5 and the train using the ph low dim dataset from robomimic  [32], we developed an agent based on behavioral cloning [32]. Robot’s actions are $\Delta$ differences. Our objective is to explain which concepts of the inputs, $C_{I}$={object, EEF position measurement, EEF quaternion measurement,gripper open width}, is responsible when the robot is taking a set of actions defined by $C_{A}=\mathrm{top~{}75\%~{}of~{}each~{}}\Delta~{}\text{action}$.
@@ IV-B Experiment 2: Lift Cube, Pick & Place, and Nut Assembly with Proprioceptive Sensors
IV-C Experiment 3: Lift Cube with Vision-Language Inputs
Setup: This experiment replicates the setup and conditions of Experiment IV-B, but proprioceptive sensors are replaced by a camera and the object information is not given to the model. The architecture is shown in Fig 6.
Figure 6: Model Architecture
Results: We evaluated $C_{I}$=(images={cube, gripper, table}, language={proper language commands, gibberish, verbs}). Interestingly, language concepts reveals that the verb in the sentence matters more than the rest of the sentence at the time it lifts the object.
Figure 7: (a) Indicates verbs in a language command matter most when it is lifting. (b) Indicates the images concepts are highly uncertain.
IV-D Experiment 4: Autonomous Driving with Vision
Setup: To simulate autonomous driving, we trained an end-to-end deep reinforcement learning policy using proximal policy optimization (PPO) [33] to steer and throttle a vehicle on Donkey Simulator. The network is a CNN followed by an MLP, trained end-to-end.
@@ IV-D Experiment 4: Autonomous Driving with Vision
Figure 8: (a) The relative importance of concepts indicates that the agent relies on the black concept ($\sim$ road) than the orange concept ($\sim$ cones). (b) Temporal change of scores. Scores for $C_{A}$ in blue.
V Limitations
We cannot provide formal guarantees that BaTCAVe will surely explain because the quality of explanations depends on the concepts we decide to test. We plan to develop automatic concept discovery methods [20, 26] tailored to robotics. Further, certain concepts are difficult to test. For instance, if we test the concept of “farther away cones” in experiment 4, the activation strength is not strong enough as only a few pixels contain the signal.
@@ VI Conclusion
We proposed a task-agnostic explainable robotics technique. We demonstrated how our method can be used to explain various robot behaviors across a variety of domains with different decision networks. The actionable insights provided by BaTCAVe helps engineers identify vulnerabilities of various components of robot training—data augmentation, fine-tuning, domain-shift analysis, verification, etc. We showed how uncertainty that BatCAVe provides helps with trusting explanations. Future work will focus on developing concept dictionaries for different robotic tasks.
@@ II-C Additional Results
Pick & place $\Delta$ X ()($\uparrow$)( ↑ ) $1\pm 0.0$ $0.54\pm 0.49$ $0.96\pm 0.16$ $0.09\pm 0.09$
$\Delta$ Y ()($\uparrow$)( ↑ ) $0.0\pm 0$ 0.0 $0.7\pm 0.41$ $0.24\pm 0.43$ $0.0\pm 0$ 0.0
$\Delta$ Z ()($\uparrow$)( ↑ ) $0.9\pm 0.17$ $0.85\pm 0.35$ $0.77\pm 0.41$ $0.91\pm 0.27$
$\Delta$ Roll ()($\uparrow$)( ↑ ) $0.44\pm 0.49$ $0.53\pm 0.49$ $0.43\pm 0.49$ $0.49\pm 0.49$
$\Delta$ Pitch ()($\uparrow$)( ↑ ) $0.86\pm 0.33$ $0.55\pm 0.49$ $0.45\pm 0.49$ $0.46\pm 0.49$
$\Delta$ Yaw ()($\uparrow$)( ↑ ) $0.84\pm 0.35$ $0.51\pm 0.49$ $0.50\pm 0.49$ $0.58\pm 0.49$
Gripper ()($\uparrow$)( ↑ ) $0.1\pm 0.0$ $0.64\pm 0.46$ $0.56\pm 0.48$ N.A*
Nut assembly $\Delta$ X ()($\uparrow$)( ↑ ) $0.64\pm 0.47$ $0.59\pm 0.49$ $0.44\pm 0.49$ $0.87\pm 0.32$
$\Delta$ Y ()($\uparrow$)( ↑ ) $0.87\pm 0.32$ $0.59\pm 0.48$ $0.46\pm 0.49$ $0.37\pm 0.48$
$\Delta$ Z ()($\uparrow$)( ↑ ) $0.005\pm 0.07$ $0.54\pm 0.49$ $0.51\pm 0.49$ $0.85\pm 0.35$
$\Delta$ Roll ()($\uparrow$)( ↑ ) $0.52\pm 0.49$ $0.48\pm 0.49$ $0.48\pm 0.49$ $0.45\pm 0.49$
$\Delta$ Pitch ()($\uparrow$)( ↑ ) $0.66\pm 0.46$ $0.51\pm 0.49$ $0.45\pm 0.49$ $0.56\pm 0.49$
$\Delta$ Yaw ()($\uparrow$)( ↑ ) $0.58\pm 0.49$ $0.44\pm 0.49$ $0.45\pm 0.49$ $0.52\pm 0.49$
Gripper ()($\uparrow$)( ↑ ) $0.1\pm 0.0$ $0.88\pm 0.31$ $0.53\pm 0.47$ N.A*
Lift $\Delta$ X ()($\uparrow$)( ↑ ) $0.84\pm 0.31$ $0.42\pm 0.40$ $0.42\pm 0.49$ $0.42\pm 0.49$
$\Delta$ Y ()($\uparrow$)( ↑ ) $0.85\pm 0.32$ $0.99\pm 0.07$ $0.56\pm 0.49$ $0.49\pm 0.40$
$\Delta$ Z ()($\uparrow$)( ↑ ) $0.98\pm 0.10$ $0.99\pm 0.04$ $0.42\pm 0.49$ $0.39\pm 0.48$
$\Delta$ Roll ()($\uparrow$)( ↑ ) $0.46\pm 0.49$ $0.41\pm 0.49$ $0.48\pm 0.48$ $0.51\pm 0.49$
$\Delta$ Pitch ()($\uparrow$)( ↑ ) $0.54\pm 0.49$ $0.28\pm 0.73$ $0.43\pm 0.47$ $0.48\pm 0.49$
$\Delta$ Yaw ()($\uparrow$)( ↑ ) $0.73\pm 0.39$ $0.28\pm 0.45$ $0.47\pm 0.49$ $0.48\pm 0.49$
Gripper ()($\uparrow$)( ↑ ) $0.95\pm 0.18$ $0.004\pm 0.06$ $0.31\pm 0.39$ N.A*
@@ IV-B Concepts
As highlighted in section V as one of the limitations, when the concepts are small, they might not provide a stong enough signal to distinguish from a random concept class. For instance, Cone Concept B and Random Concept in Fig. 25 are largely similar.

### v2

@@ IV-C Interpreting the Trustworthiness of Explanations
The thresholds for probability should be decided by the practitioners depending on how much risk they are willing to take. For instance, if an engineer is using BaTCAVe to debug a manipulator used in an assembly line, the threshold can be selected leniently as the stakes might be relatively low. In such cases, we can obtain more valid explanations. In contrast, if a legislative body is using BaTCAVe to approve a new autonomous vehicle, then the thresholds should be strict. If case 1 violates, we should try new concepts to obtain a better accuracy. If case 2 violates, we can still use explanations but they might not always be the best explanations. By obtaining the mean score, we can obtain an average explanation.
V Experiments
We consider three types of algorithms for training robot decision-making: binary supervised learning, behavioral cloning (BC), and proximal policy optimization (PPO). The experimental setup, results, and findings are detailed below.
TABLE I: BaTCAVe scores $(\uparrow)$ across different tasks, quantifying the relevance of each action concept to the task, along with uncertainty estimates. Each score measures the impact of input concepts such as the object’s features, end-effector position (eef_pos), end-effector orientation (eef_quat in quaternion format), and gripper on task performance. (*Not Applicable)
Tasks Action Concepts Input Concepts Conclusions
object eef_pos eef_quat gripper
Pick & Place[Uncaptioned image] $\Delta$ X $\mathbf{1.00\pm 0.00}$ $0.54\pm 0.49$ $\mathbf{0.96\pm 0.16}$ $0.09\pm 0.09$ Object features are most important, with $\Delta X$ (1.00) and $\Delta Z$ (0.90) scoring highest. End-effector orientation ($0.96$) aids alignment and gripper is relevant for $\Delta Z$ ($0.91$).
$\Delta$ Y $0.00\pm 0.00$ $0.7\pm 0.41$ $24\pm 0.43 0.0$$00\pm 0.00$
$\Delta$ Z $\mathbf{0.90\pm 0.17}$ $0.85\pm 0.35$ $0.77\pm 0.41$ $\mathbf{0.91\pm 0.27}$
$\Delta$ Roll $0.44\pm 0.49$ $0.53\pm 0.49$ $0.43\pm 0.49$ $0.49\pm 0.49$
$\Delta$ Pitch $0.86\pm 0.33$ $0.55\pm 0.49$ $0.45\pm 0.49$ $0.46\pm 0.49$
$\Delta$ Yaw $0.84\pm 0.35$ $0.51\pm 0.49$ $0.50\pm 0.49$ $0.58\pm 0.49$
Gripper $0.10\pm 0.00$ $0.64\pm 0.46$ $0.56\pm 0.48$ N.A*
Lift[Uncaptioned image] $\Delta$ X $0.84\pm 0.31$ $0.42\pm 0.40$ $0.42\pm 0.49$ $0.42\pm 0.49$ End-effector positioning is critical, with $\Delta Y$ ($0.99$) and $\Delta Z$ ($0.99$) dominating. The gripper’s role ($0.95$) suggests a strong grasp is essential.
$\Delta$ Y $0.85\pm 0.32$ $\mathbf{0.99\pm 0.07}$ $0.56\pm 0.49$ $0.49\pm 0.40$
@@ V Experiments
$\Delta$ Roll $0.46\pm 0.49$ $0.41\pm 0.49$ $0.48\pm 0.48$ $0.51\pm 0.49$
$\Delta$ Pitch $0.54\pm 0.49$ $0.28\pm 0.73$ $0.43\pm 0.47$ $0.48\pm 0.49$
$\Delta$ Yaw $0.73\pm 0.39$ $0.28\pm 0.45$ $0.47\pm 0.49$ $0.48\pm 0.49$
Gripper $\mathbf{0.95\pm 0.18}$ $0.004\pm 0.06$ $0.31\pm 0.39$ N.A*
Nut Assembly[Uncaptioned image] $\Delta$ X $0.64\pm 0.47$ $0.59\pm 0.49$ $0.44\pm 0.49$ $\mathbf{0.87\pm 0.32}$ Gripper control is key, influencing $\Delta X$ ($0.87$) and $\Delta Z$ ($0.85$). Object and end-effector positioning contribute moderately, indicating precise handling.
$\Delta$ Y $\mathbf{0.87\pm 0.32}$ $0.59\pm 0.48$ $0.46\pm 0.49$ $0.37\pm 0.48$
$\Delta$ Z $0.01\pm 0.07$ $0.54\pm 0.49$ $0.51\pm 0.49$ $\mathbf{0.85\pm 0.35}$
$\Delta$ Roll $0.52\pm 0.49$ $0.48\pm 0.49$ $0.48\pm 0.49$ $0.45\pm 0.49$
$\Delta$ Pitch $0.66\pm 0.46$ $0.51\pm 0.49$ $0.45\pm 0.49$ $0.56\pm 0.49$
$\Delta$ Yaw $0.58\pm 0.49$ $0.44\pm 0.49$ $0.45\pm 0.49$ $0.52\pm 0.49$
Gripper $0.10\pm 0.00$ $0.88\pm 0.31$ $0.53\pm 0.47$ N.A*
V-A Experiment 1: Vision-based Mobile Robot Navigation
Setup: As a real-world platform, we used a JetBot with a NVIDIA Jetson Nano (Fig 5). It uses a pretrained AlexNet [27] with the last layer replaced with two nodes. Using 175 “obstacle” and 175 “free” images, we fine-tune the DNN so that the robot can learn to avoid obstacles. With the objective of avoiding orange obstacles, we fine-tuned the full DNN with images of an orange box, shown in Fig. 5, until we achieve a validation accuracy of 100%. We analyze how the decision of the network is affected by different data pre-processing and fine-tuning techniques for $C_{A}=\{\text{avoid obstacle}\}$ and $C_{I}=\{\text{orangeness, darkness, distance}\}$.
Results:
Q1. Are explanations useful to the engineers? We surveyed 20 engineering students who have at least 1 year of experience training DNNs to get their opinion on what the DNN has learned just by reading our description. We described them the architecture and showed fine-tuning images without telling them that our objective is avoiding orange obstacles.
First, we asked participants to describe what attributes the DNN should have learned and then we gave them a list of potential concepts to narrow down their choices. As shown in Figs. 4(a), many participants (35%) speculated that the model would have learned to distinguish orange from the rest. In contrast, showcasing the BaTCAVe’s ability to provide true explanations that engineers might not even think of, as shown in Fig. 4(b), BaTCAVe revealed that the model had learned to distinguish dark from light objects (or the value or brightness in HSV scale). The orange box is merely a shade of the broader “dark” concept. Thus, we conclude that BaTCAVe is capable of providing explanations of inner understanding of the neural network, which, without explanations, is not intuitive to engineers.
Q2. Are BaTCAVe explanations correct? Similar to C-deletion in XAI literature [19, 20, 21], we verified that the darkness is an important concept by gradually varying the brightness of the orange object, as shown in Fig. 4(f). Based on Fig. 4, BaTCAVe made the following explanations:
  1. 1.
@@ Results:
  3. 3.
    As shown in Fig 4(c), the importance of the dark concept remains consistent across different data augmentation methods—adding color jitter (CJ) and/or image rotation (R)—for LFT while it varies for FFT.
Since XAI methods are hard to quantitatively benchmark, similar to other work, Fig. 4(e) qualitatively compares why BaTCAVe is a better choice than feature attribution methods for robot learning.
V-B Experiment 2: Lift Cube, Pick & Place, and Nut Assembly with Proprioceptive Sensors
Setup: We use a Panda, a 6-DOF robotic arm with a gripper, to complete the three tasks shown in Table I on RoboSuite [28]. The setup includes various proprioceptive sensors to monitor the arm’s movements and positions. By using the DNN shown in Fig 6a and trained on the low-dimensional dataset from robomimic  [29], we developed an agent based on behavioral cloning [29]. Robot’s actions are $\Delta$ differences. Our objective is to explain which concepts of the inputs, $C_{I}$={object, EEF position measurement, EEF quaternion measurement,gripper open width}, is responsible when the robot is taking a set of actions defined by $C_{A}=\mathrm{top~75\%~of~each~}\Delta~\text{action}$.
@@ V-B Experiment 2: Lift Cube, Pick & Place, and Nut Assembly with Proprioceptive Sensors
V-C Experiment 3: Lift Cube with Vision-Language Inputs
Setup: As a complex SOTA task, this experiment replicates the setup and conditions of Experiment V-B, but proprioceptive sensors are replaced by a high-resolution RGB images ($\times224$$\times$3) from the camera and language instructions describing the task. The model was train for 100,000 epochs on 300 demonstrations with a batch size of 64 using a Huber loss function and Adam optimizer. The architecture of the network is shown in Fig 7.
Figure 7: Model architecture for robot control used in Sec V-C, consisting of a ResNet-based Visual Encoder, a Visual Narrower for dimensionality reduction, a CLIP-based Task Encoder. The final linear layer outputs a 360-dimensional action vector, mapping the robot’s motion over 36 timesteps.
Results: We evaluated $C_{I}$=(images={cube, gripper, table}, language={proper language commands, gibberish, verbs}). Interestingly, language concepts reveals that the verb in the sentence matters more than the rest of the sentence at the time it lifts the object as seen in Fig 8(a) where verbs have high score with low uncertainty towards the end of the episode for $C_{A}=X$.
Figure 8: (a) Indicates verbs in a language command matter most when it is lifting. (b) Indicates the images concepts are highly uncertainy.
@@ V-C Experiment 3: Lift Cube with Vision-Language Inputs
Figure 9: (a) The relative importance of concepts indicates that the agent relies on the black concept ($\sim$ road) than the orange concept ($\sim$ cones). (b) Temporal change of scores.
V-D Experiment 4: Autonomous Driving with Vision
Setup: To simulate autonomous driving, we trained an end-to-end deep reinforcement learning policy using proximal policy optimization (PPO) [30] to steer and throttle a vehicle on Donkey Simulator. The network is a CNN policy followed by an MLP, which was trained for a total of 100,000 timesteps with learning rate of 0.0003, a rollout of 2048 steps per update, and a batch size of 64.
@@ VI Conclusion
We proposed a task-agnostic explainable robotics technique. We demonstrated how our method can be used to explain various robot behaviors across a variety of domains with different decision networks. The actionable insights provided by BaTCAVe helps engineers identify vulnerabilities in various components of robot training—data augmentation, fine-tuning, domain-shift analysis, verification, etc. We show how the uncertainty provided by BatCAVe helps in distinguishing between reliable and ambiguous explanations, which is crucial for high-stakes applications. Future work will focus on explaining sequential decisions rather than specific behaviors and developing methods to amplify activations that have limited signal (e.g., only a few pixels cover “farther away cones” in experiment 4).
@@ II-C Additional Results
Pick & place $\Delta$ X $(\uparrow)$ $1\pm 0.0$ $0.54\pm 0.49$ $96\pm 0.16 0.0$$09\pm 0.09$
$\Delta$ Y $(\uparrow)$ $0.0\pm 0.0$ $0.7\pm 0.41$ $24\pm 0.43 0.0$$0\pm 0.0$
$\Delta$ Z $(\uparrow)$ $0.9\pm 0.17$ $0.85\pm 0.35$ $0.77\pm 0.41$ $0.91\pm 0.27$
$\Delta$ Roll $(\uparrow)$ $0.44\pm 0.49$ $0.53\pm 0.49$ $0.43\pm 0.49$ $0.49\pm 0.49$
$\Delta$ Pitch $(\uparrow)$ $0.86\pm 0.33$ $0.55\pm 0.49$ $0.45\pm 0.49$ $0.46\pm 0.49$
$\Delta$ Yaw $(\uparrow)$ $0.84\pm 0.35$ $0.51\pm 0.49$ $0.50\pm 0.49$ $0.58\pm 0.49$
Gripper $(\uparrow)$ $0.1\pm 0.0$ $0.64\pm 0.46$ $0.56\pm 0.48$ N.A*
Nut assembly $\Delta$ X $(\uparrow)$ $0.64\pm 0.47$ $0.59\pm 0.49$ $0.44\pm 0.49$ $0.87\pm 0.32$
$\Delta$ Y $(\uparrow)$ $0.87\pm 0.32$ $0.59\pm 0.48$ $0.46\pm 0.49$ $0.37\pm 0.48$
$\Delta$ Z $(\uparrow)$ $0.005\pm 0.07$ $0.54\pm 0.49$ $0.51\pm 0.49$ $0.85\pm 0.35$
$\Delta$ Roll $(\uparrow)$ $0.52\pm 0.49$ $0.48\pm 0.49$ $0.48\pm 0.49$ $0.45\pm 0.49$
$\Delta$ Pitch $(\uparrow)$ $0.66\pm 0.46$ $0.51\pm 0.49$ $0.45\pm 0.49$ $0.56\pm 0.49$
$\Delta$ Yaw $(\uparrow)$ $0.58\pm 0.49$ $0.44\pm 0.49$ $0.45\pm 0.49$ $0.52\pm 0.49$
Gripper $(\uparrow)$ $0.1\pm 0.0$ $0.88\pm 0.31$ $0.53\pm 0.47$ N.A*
Lift $\Delta$ X $(\uparrow)$ $0.84\pm 0.31$ $0.42\pm 0.40$ $0.42\pm 0.49$ $0.42\pm 0.49$
$\Delta$ Y $(\uparrow)$ $0.85\pm 0.32$ $0.99\pm 0.07$ $0.56\pm 0.49$ $0.49\pm 0.40$
$\Delta$ Z $(\uparrow)$ $8\pm 0.10 0.99±0$$9\pm 0.04$ $0.42\pm 0.49$ $0.39\pm 0.48$
$\Delta$ Roll $(\uparrow)$ $0.46\pm 0.49$ $0.41\pm 0.49$ $0.48\pm 0.48$ $0.51\pm 0.49$
$\Delta$ Pitch $(\uparrow)$ $0.54\pm 0.49$ $0.28\pm 0.73$ $0.43\pm 0.47$ $0.48\pm 0.49$
$\Delta$ Yaw $(\uparrow)$ $0.73\pm 0.39$ $0.28\pm 0.45$ $0.47\pm 0.49$ $0.48\pm 0.49$
Gripper $(\uparrow)$ $95\pm 0.18 0.0$$004\pm 0.06$ $0.31\pm 0.39$ N.A*
@@ IV-B Concepts
One of the limitations are when the concepts are small, they might not provide a strong enough signal to distinguish from a random concept class. For instance, Cone Concept B and Random Concept in Fig. 26 are largely similar.